QUESTIONS 6 AND 7 ARE BASED ON THE FOLLOWING INFORMATION.

A Check found that sources of fake news on Facebook are mostly ghost profiles. Suppose
that 20% of profiles on Facebook are ghost profiles. Suppose further that we randomly select 20
Facebook profiles and check whether or not they are ghost profiles.

Question 6

What is the probability that at least 11 profiles are ghost profiles?

(1) 0.0026
(2) 0.0020
(3) 0.0002
(4) 0.0139
(5) None of the above.

Question 7

What is the value of the mean and the standard deviation?

(1) µ = 4 and σ = 3.2
(2) µ = 4 and σ = 1.79
(3) µ = 20 and σ = 1.79
(4) µ = 20 and σ = 3.2
(5) Need more information to determine both values.
1
Expert's answer
2020-05-11T17:42:27-0400

6) (5)

µ=n*p=20*0.2=4

σ="\sqrt{npq}" ="\sqrt{20*0.8*0.2}" =1.79

P"(x \geq11)"=1-P(x<11)

z=(x-µ)/σ=(11-4)/1.79=3.91

using table P(x<11)=0.99995

P(x>=11)=1-P(x<11)=0.00005

7) (2)

µ=n*p=20*0.2=4

σ="\sqrt{npq}" = "\sqrt{20*0.8*0.2}" =1.79